# TreeDGS: Aerial Gaussian Splatting for Distant DBH Measurement

### v2

@@ 3.3 Reconstruction with Gaussian Splats
Given calibrated cameras $\{\mathbf{P}_{i}\}$ and an MVS-densified initialization $\mathcal{X}_{\text{dense}}$ (Sec. 3.2), we optimize a 3D Gaussian Splatting (3DGS) scene representation. We adopt RaDe-GS [50], which builds on the real-time 3DGS renderer [22] while explicitly rasterizing depth in a way that is also useful for our subsequent surface-consistent point sampling.
Gaussian parameterization.
The scene is represented as a set of anisotropic 3D Gaussians $G_{k}=(\boldsymbol{\mu}_{k},\mathbf{s}_{k},\mathbf{R}_{k},\alpha_{k},\mathbf{c}_{k})$, where $\boldsymbol{\mu}_{k}\in\mathbb{R}^{3}$ is the Gaussian center, $\mathbf{s}_{k}\in\mathbb{R}^{3}_{+}$ are axis-aligned scales in the local frame, $\mathbf{R}_{k}\in SO(3)$ is the per-Gaussian orientation, $\alpha_{k}\in[0,1]$ is the learned opacity, and $\mathbf{c}_{k}$ denotes appearance parameters (e.g., spherical-harmonic features as in [22]). We initialize $\{\boldsymbol{\mu}_{k}\}$ from $\mathcal{X}_{\text{dense}}$ and optimize all parameters following RaDe-GS [50].
RaDe-GS depth-plane formulation.
A key difference from vanilla 3DGS is that RaDe-GS associates each Gaussian and view with a local ray-distance plane in screen space. For a view $v$, let $\mathbf{u}_{k,v}\in\mathbb{R}^{2}$ be the projected mean of Gaussian $k$, and let $t_{k,v}$ be its ray-distance (range) in that view (i.e., $\|\mathbf{R}_{v}\boldsymbol{\mu}_{k}+\mathbf{t}_{v}\|_{2}$ under perspective projection). RaDe-GS additionally provides a 2D slope $\mathbf{g}_{k,v}\in\mathbb{R}^{2}$ , so that the Gaussian’s predicted ray-distance at a nearby pixel $\mathbf{u}$ is approximated by
@@
Adaptive training with RLGS.
We integrate RLGS [28] as an online controller that adapts selected training hyper-parameters based on observed optimization dynamics. In our setting, this is used to stabilize opacity and geometry quality for the downstream DBH pipeline. After convergence, the optimized Gaussian field is denoted $\mathcal{G}$ and serves as the source representation for dense point sampling.
@@ 3.4 Surface Sampling with Opacity
Our downstream stages (3D trunk segmentation and DBH fitting) require a point cloud that is (i) dense enough to represent thin stems at stand-off distance and (ii) geometrically trustworthy under heavy occlusion from foliage and understory. This is not automatic from a 3D Gaussian Splatting (3DGS) reconstruction: the optimized field $\mathcal{G}$ is a collection of overlapping volumetric primitives whose parameters are trained through front-to-back alpha compositing rather than an explicit surface loss [22, 50]. As a result, two seemingly straightforward exports are unreliable in our UAV forest regime. First, exporting one point per Gaussian (the means) produces a fragmented and extremely sparse cloud (Sec. 4, Fig. 10), which is insufficient for robust 3D segmentation and measurement. Second, back-projecting rendered depth maps into 3D and fusing them via multi-view depth-consistency preserves view-consistent sheet-like artifacts from semi-transparent vegetation while eroding already-scarce trunk evidence, especially when stems are only a few pixels wide or intermittently visible (Fig. 6). Therefore, we sample densely in 3D, but we accept and score samples using tests that are consistent with the same compositing model that governs 3DGS rendering and optimization.
Our goal is to convert the optimized Gaussian field $\mathcal{G}$ into a point cloud that is dense enough for trunk segmentation and DBH fitting. For each Gaussian $G_{i}=(\boldsymbol{\mu}_{i},\mathbf{s}_{i},\mathbf{R}_{i},\alpha_{i})$, we draw candidate offsets $\boldsymbol{\xi}_{ij}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{3})$ in the local frame and map them to world coordinates. To focus sampling on reliable (high-opacity) regions, we apply Bernoulli thinning controlled by $\alpha_{i}$:
@@
We keep $\mathbf{p}_{ij}$ if $b_{ij}=1$; in expectation each Gaussian contributes $\alpha_{i}M$ points when drawing $M$ candidates.
In 3DGS, solidity is not defined by an explicit surface. It is defined implicitly by how Gaussians accumulate opacity under the renderer’s front-to-back alpha compositing. Therefore, any surface sampling rule that aims to match what the model learned should be expressed using the same compositing mechanism. For a view $v$ and pixel $\mathbf{u}$, the renderer composites Gaussians front-to-back using transmittance tracking. If $\alpha_{k,v}(\mathbf{u})$ is the per-Gaussian alpha contribution at $\mathbf{u}$, we define $T_{0}(\mathbf{u})=1$ and update
@@
The per-pixel accumulated alpha mask $A_{v}(\mathbf{u})$ tells us whether a ray intersects reconstructed mass somewhere along the ray, but it does not tell us whether a particular 3D sample $\mathbf{p}$ lies on the visible surface, behind it, or in front of it (all can project to the same pixel). For surface extraction, we need a depth-aware test that measures how much opacity accumulates in front of $\mathbf{p}$ along the viewing ray, using the same compositing model that governs training. RaDe-GS provides exactly this query by integrating point-wise alphas with a depth-plane formulation.
Beyond per-pixel alpha, RaDe-GS provides an integration query at an arbitrary 3D point $\mathbf{p}$: it returns the projected pixel coordinate $\mathbf{u}_{v}(\mathbf{p})$ and a cumulative opacity value $\tilde{\alpha}_{v}(\mathbf{p})$ that measures how much opacity is accumulated in front of $\mathbf{p}$ along the viewing ray. Let $r_{v}(\mathbf{p})=\|\mathbf{R}_{v}\mathbf{p}+\mathbf{t}_{v}\|_{2}$ denote the point ray-distance in view $v$. For a Gaussian $k$, RaDe-GS evaluates a ray-space quadratic form using a precomputed inverse covariance over the 3D offset
@@
The $\min(\cdot)$ depth clamping in Eq. (6) is critical for surface consistency: once $\mathbf{p}$ lies behind the locally rasterized surface depth $d_{k,v}(\cdot)$, the depth residual term saturates rather than increases with $r_{v}(\mathbf{p})$, preventing spurious “volumetric” accumulation behind the surface. RaDe-GS composites these point-wise alphas using the same front-to-back transmittance as in Eq. (5), yielding the cumulative opacity
@@
Since DBH fitting is highly sensitive to even small amounts of geometric contamination, we aggregate evidence across the calibrated views and attach a per-point reliability that downstream steps can threshold or use as a weight. Concretely, we (i) only allow a view to contribute if the point projects to a foreground pixel in that view, and (ii) summarize the point-wise opacity integrals across contributing views conservatively. Let $\mathcal{P}=\{\mathbf{p}_{ij}\mid b_{ij}=1\}$ be the candidate samples retained after the opacity-guided thinning in Eq. (4). For each $\mathbf{p}\in\mathcal{P}$, we evaluate its support over the calibrated views; a view $v$ contributes only if $\mathbf{p}$ projects inside the image and lands on a non-background pixel in the rendered accumulated alpha mask $A_{v}$. Specifically, with $\mathbf{u}_{v}(\mathbf{p})$ the projection of $\mathbf{p}$, we define
$m_{v}(\mathbf{p})=\mathbf{1}\!\left(A_{v}(\mathbf{u}_{v}(\mathbf{p}))>\tau_{\text{mask}}\right).$ (9)
@@
Here $\bar{\alpha}(\mathbf{p})$ is a conservative multi-view opacity estimate, and $w(\mathbf{p})$ counts the number of views in which $\mathbf{p}$ projects onto a foreground pixel, defined as a pixel whose rendered alpha exceeds $\tau$. In practice, for surface extraction we keep points with $\bar{\alpha}(\mathbf{p})>\tau$ and $w(\mathbf{p})>0$ (we use $\tau=0.5$); for dense export we set $\tau=0$ and keep all points while storing $(\bar{\alpha},w)$ for downstream weighting.
@@ 3.5 Semantic Trunk Extraction
The sampled point cloud from Section 3.4 contains trunks mixed with foliage, understory vegetation, and occasional floating samples caused by Gaussians with nonzero opacity in free space. To isolate stem geometry for DBH measurement, we used ForestFormer3D [48], a 3D semantic segmentation model, to predict a class label for each sampled point and to assign per-tree instance IDs when available. In our pipeline, the per-point opacity reliability $\bar{\alpha}$ is kept as an auxiliary output feature alongside $(x,y,z)$. We retain only points predicted as trunk to form a trunk-only cloud for each tree instance $t$, denoted $\mathcal{T}_{t}=\{(\mathbf{x}_{k},\bar{\alpha}_{k})\}$. This separation step reduces contamination from branches and understory, while the retained opacity values enable reliability-weighted fitting downstream. Then, the trunk-only instances are passed to the DBH measurement stage.
@@ 3.6 Opacity-Weighted DBH Measurement
DBH is defined relative to local ground height. For each tree instance, we estimate a ground elevation $z_{g}$ by querying a digital terrain model (DTM) at the tree location (e.g., nearest-neighbor lookup at the mean trunk $(x,y)$), and convert all points to height-above-ground:
@@
All subsequent slicing and DBH evaluation are performed in the $(x,y)$ plane as a function of $h$.
We construct a sequence of slice centers $\{h_{s}\}$ with spacing $\Delta z$ and thickness $H$ (a slab of height $H$), and collect 2D points in each slab:
@@
Slices with fewer than a small minimum number of points (e.g., <5<5) are discarded. This step produces a set of candidate diameters along the lower stem.
Unlike classical TLS stem fitting (where points often lie near the circumference), our sampling in Section 3.4 draws points from Gaussian volumes; therefore, many slice points legitimately fall inside the trunk cross-section. For this reason we treat inliers as belonging to a filled disk rather than a narrow ring around a circle. This choice makes the estimator consistent with volumetric sampling and substantially more stable under partial visibility.
Opacity-weighted solid-circle RANSAC in each slice.
For slice $s$, we have 2D points $\{(\mathbf{q}_{k},w_{k})\}_{k=1}^{n}$ with $\mathbf{q}_{k}=(x_{k},y_{k})^{\top}$ and opacity weights $w_{k}=\bar{\alpha}_{k}$. Each RANSAC hypothesis performs the following steps:
@@
(ii) Circle fit from three points: Fit the circle through the sampled points using the algebraic form $x^{2}+y^{2}+ax+by+d=0$. We solve for $(a,b,d)$ after subtracting the slice mean (for numerical stability), then recover $\mathbf{c}=(-\tfrac{a}{2},-\tfrac{b}{2})^{\top}$ and $r^{2}=\lVert\mathbf{c}\rVert_{2}^{2}-d$ (adding the mean back to $\mathbf{c}$).
@@
Height-wise taper regression and DBH prediction.
Even after per-slice robust fitting, some slices can be corrupted by branch attachments, residual foliage, or incomplete sampling. Therefore, we fit a simple taper model on the lower stem using RANSAC with a negative-slope prior:
$\hat{d}(h)=\beta_{0}+\beta_{1}h,\qquad\beta_{1}<0,$ (15)
where RANSAC discards outlier slices and the slope constraint enforces physically plausible taper (diameter should not increase with height over short stem segments). In practice, we fit Eq. (15) over progressively larger height ranges starting near the ground (to avoid upper-canopy contamination) until a stable inlier set is found. Finally, DBH is reported as
@@
Role of opacity in robustness.
The opacity weights $w_{k}=\bar{\alpha}_{k}$ originate from the multi-view rendering consistency test in Section 3.4. Points that are only weakly supported (e.g., floaters, thin vegetation, or ambiguous geometry) tend to have small $\bar{\alpha}$ and therefore (i) are sampled less often during hypothesis generation and (ii) contribute little to the inlier score in Eq. (14). This coupling between rendering-consistency and geometric fitting is the central mechanism behind our opacity-weighted measurement mode.
Table 2: Per-plot DBH accuracy and robustness. RMSE/ MAE/ME are in cm and RRMSE is in %. SR reports successful estimates over total trees. Bold indicates the best within each plot. TreeDGS with opacity-weighted circle fitting attains the best RMSE in 7/10 plots and the best aggregated performance.
Plot Method RMSE RRMSE MAE ME SR
(cm) (%) (cm) (cm) (#/#)
All LiDAR (Cyl. fit.) 7.91 28.85 5.04 1.64 187/190
LiDAR (Cir. nw. fit.) 9.36 33.96 8.23 8.11 183/190
LiDAR (Cir. w. fit.) 10.99 39.64 9.42 7.16 179/190
TreeDGS (Cyl. fit.) 13.29 48.69 10.25 -6.46 189/190
TreeDGS (Cir. nw. fit.) 9.61 35.22 6.97 5.79 189/190
TreeDGS (Cir. w. fit.) 4.79 17.54 3.70 -0.38 189/190
1 LiDAR (Cyl. fit.) 6.09 19.53 4.68 -1.66 19/19
LiDAR (Cir. nw. fit.) 6.58 21.10 5.62 4.60 19/19
LiDAR (Cir. w. fit.) 6.25 20.13 5.50 5.01 18/19
TreeDGS (Cyl. fit.) 11.69 37.47 10.99 -10.99 19/19
TreeDGS (Cir. nw. fit.) 7.58 24.29 5.91 2.63 19/19
TreeDGS (Cir. w. fit.) 4.94 15.83 3.72 -2.04 19/19
2 LiDAR (Cyl. fit.) 4.73 16.63 4.04 1.86 22/22
LiDAR (Cir. nw. fit.) 8.87 31.20 7.68 7.68 22/22
LiDAR (Cir. w. fit.) 8.47 29.80 7.28 4.01 22/22
TreeDGS (Cyl. fit.) 13.43 47.21 9.84 -5.49 22/22
TreeDGS (Cir. nw. fit.) 8.43 29.65 7.26 7.09 22/22
TreeDGS (Cir. w. fit.) 5.07 17.82 3.18 1.88 22/22
3 LiDAR (Cyl. fit.) 17.22 66.04 9.85 7.77 20/20
LiDAR (Cir. nw. fit.) 10.26 37.93 9.04 9.04 19/20
LiDAR (Cir. w. fit.) 14.81 56.79 12.09 10.60 20/20
TreeDGS (Cyl. fit.) 10.49 40.23 8.77 -7.35 20/20
TreeDGS (Cir. nw. fit.) 16.39 62.83 9.47 8.15 20/20
TreeDGS (Cir. w. fit.) 5.93 22.73 4.35 -0.10 20/20
4 LiDAR (Cyl. fit.) 6.24 25.47 5.15 4.12 16/16
LiDAR (Cir. nw. fit.) 11.20 45.69 10.45 10.45 16/16
LiDAR (Cir. w. fit.) 9.61 38.40 7.66 4.79 15/16
TreeDGS (Cyl. fit.) 25.24 102.93 16.57 3.84 16/16
TreeDGS (Cir. nw. fit.) 11.46 46.73 7.59 4.95 16/16
TreeDGS (Cir. w. fit.) 5.74 23.40 4.53 -0.62 16/16
5 LiDAR (Cyl. fit.) 5.27 18.60 3.83 1.36 21/21
LiDAR (Cir. nw. fit.) 8.42 29.69 7.92 7.92 20/21
LiDAR (Cir. w. fit.) 13.62 47.99 12.86 12.86 20/21
TreeDGS (Cyl. fit.) 12.39 43.69 11.01 -7.98 21/21
TreeDGS (Cir. nw. fit.) 5.84 20.60 4.64 3.57 21/21
TreeDGS (Cir. w. fit.) 4.38 15.46 4.01 -2.10 21/21
6 LiDAR (Cyl. fit.) 6.83 25.38 4.67 1.24 22/22
LiDAR (Cir. nw. fit.) 8.91 33.09 7.75 7.59 22/22
LiDAR (Cir. w. fit.) 10.99 40.83 9.69 8.56 22/22
TreeDGS (Cyl. fit.) 8.64 32.10 7.91 -7.45 22/22
TreeDGS (Cir. nw. fit.) 8.91 33.10 8.23 8.23 22/22
TreeDGS (Cir. w. fit.) 3.15 11.72 2.48 0.91 22/22
7 LiDAR (Cyl. fit.) 5.06 23.52 4.18 0.04 15/17
LiDAR (Cir. nw. fit.) 12.15 54.45 10.76 10.76 14/17
LiDAR (Cir. w. fit.) 9.31 39.24 7.63 4.06 12/17
TreeDGS (Cyl. fit.) 13.66 65.77 8.15 -0.28 17/17
TreeDGS (Cir. nw. fit.) 12.01 57.81 10.57 9.68 17/17
TreeDGS (Cir. w. fit.) 5.91 28.44 5.03 3.27 17/17
8 LiDAR (Cyl. fit.) 7.78 26.68 6.22 2.86 16/16
LiDAR (Cir. nw. fit.) 10.06 34.52 8.42 8.42 16/16
LiDAR (Cir. w. fit.) 12.19 40.37 11.37 7.55 15/16
TreeDGS (Cyl. fit.) 11.83 40.59 10.97 -7.64 16/16
TreeDGS (Cir. nw. fit.) 5.42 18.59 3.96 3.83 16/16
TreeDGS (Cir. w. fit.) 3.68 12.63 3.14 -1.69 16/16
9 LiDAR (Cyl. fit.) 6.18 23.13 4.20 -0.38 16/16
LiDAR (Cir. nw. fit.) 10.03 38.25 9.48 9.48 15/16
LiDAR (Cir. w. fit.) 10.61 40.45 9.99 7.36 15/16
TreeDGS (Cyl. fit.) 8.73 32.66 7.73 -7.73 16/16
TreeDGS (Cir. nw. fit.) 8.50 31.81 6.85 5.17 16/16
TreeDGS (Cir. w. fit.) 3.93 14.72 3.08 -1.26 16/16
10 LiDAR (Cyl. fit.) 4.34 14.71 3.67 -0.97 20/21
LiDAR (Cir. nw. fit.) 7.51 25.42 6.76 6.76 20/21
LiDAR (Cir. w. fit.) 10.58 35.82 9.42 5.03 20/21
TreeDGS (Cyl. fit.) 12.07 40.88 11.40 -11.15 20/21
TreeDGS (Cir. nw. fit.) 6.28 21.27 5.15 4.10 20/21
TreeDGS (Cir. w. fit.) 4.42 14.95 3.73 -2.35 20/21

### v3

@@ 3.3 Reconstruction with Gaussian Splats
Given calibrated cameras $\{\mathbf{P}_{i}\}$ and an MVS-densified initialization $\mathcal{X}_{\text{dense}}$ from Section 3.2, we optimize a 3D Gaussian Splatting (3DGS) scene representation. We adopt RaDe-GS [52], which builds on the real-time 3DGS renderer [23] while explicitly rasterizing depth in a way that is also useful for our subsequent surface-consistent point sampling.
Gaussian parameterization. The scene is represented as a set of anisotropic 3D Gaussians $G_{k}=(\boldsymbol{\mu}_{k},\mathbf{s}_{k},\mathbf{R}_{k},\alpha_{k},\mathbf{c}_{k})$, where $\boldsymbol{\mu}_{k}\in\mathbb{R}^{3}$ is the Gaussian center, $\mathbf{s}_{k}\in\mathbb{R}^{3}_{+}$ are axis-aligned scales in the local frame, $\mathbf{R}_{k}\in SO(3)$ is the per-Gaussian orientation, $\alpha_{k}\in[0,1]$ is the opacity, and $\mathbf{c}_{k}$ denotes appearance parameters (e.g., spherical-harmonic features as in [23]). We initialize $\{\boldsymbol{\mu}_{k}\}$ from $\mathcal{X}_{\text{dense}}$ and optimize all parameters following RaDe-GS [52]. Note that RaDe-GS interleaves photometric optimization with density control (splitting and pruning). Therefore, the Gaussian set can be progressively densified and refined over time, improving coverage in regions that are only weakly supported by the MVS initialization.
RaDe-GS depth-plane formulation. A key difference from vanilla 3DGS is that RaDe-GS associates each Gaussian and view with a local ray-distance plane in screen space. For a view $v$, let $\mathbf{u}_{k,v}\in\mathbb{R}^{2}$ be the projected mean of Gaussian $k$, and let $t_{k,v}$ be its ray-distance (range) in that view (i.e., $\|\mathbf{R}_{v}\boldsymbol{\mu}_{k}+\mathbf{t}_{v}\|_{2}$ under perspective projection). RaDe-GS additionally provides a 2D slope $\mathbf{g}_{k,v}\in\mathbb{R}^{2}$, so that the Gaussian’s predicted ray-distance at a nearby pixel $\mathbf{u}$ is approximated by
@@
Adaptive training with RLGS. We integrate RLGS [29] as an online controller that adapts selected training hyper-parameters based on observed optimization dynamics. In our setting, this is primarily used to stabilize opacity and geometry quality for the downstream DBH pipeline. After convergence, the optimized Gaussian field is denoted by $\mathcal{G}$ and serves as the source representation for dense point sampling.
@@ 3.4 Surface Sampling with Opacity
Our downstream stages (3D trunk segmentation and DBH fitting) require a point cloud that is (i) dense enough to represent thin stems at stand-off distance and (ii) geometrically trustworthy under heavy occlusion from foliage and understory. This is not automatic from a 3D Gaussian Splatting (3DGS) reconstruction: the optimized field $\mathcal{G}$ is a collection of overlapping volumetric primitives whose parameters are trained through front-to-back alpha compositing rather than an explicit surface loss  [23, 52]. As a result, two seemingly straightforward exports are unreliable in our UAV forest regime. First, exporting one point per Gaussian (the means) produces a fragmented and extremely sparse cloud
which is insufficient for robust 3D segmentation and measurement. Second, back-projecting rendered depth maps into 3D and fusing them via multi-view depth-consistency preserves view-consistent sheet-like artifacts from semi-transparent vegetation while eroding already-scarce trunk evidence, especially when stems are only a few pixels wide or intermittently visible (see Figure 7). Therefore, we sample densely in 3D, but we accept and score samples using tests that are consistent with the same compositing model that governs 3DGS rendering and optimization.
Figure 7: Comparison of back-projection points with multi-view depth consistency and the proposed sample points. (a) Depth-fusion baseline: rendered depth points are back-projected into 3D and kept only if they pass multi-view depth-consistency checks. This produces large sheet-like artifacts near the ground/understory and incomplete stem surfaces (red dashed boxes), despite appearing consistent across views. (b) Our opacity-guided surface sampling with depth-aware point-wise compositing give a stem geometry suitable for DBH estimation. Blue indicates lower elevations near ground, and red/orange indicates higher canopy points.
Opacity-guided stochastic sampling. Our goal is to convert the optimized Gaussian field $\mathcal{G}$ into a point cloud that is dense enough for trunk segmentation and DBH fitting. For each Gaussian $G_{i}=(\boldsymbol{\mu}_{i},\mathbf{s}_{i},\mathbf{R}_{i},\alpha_{i})$, we draw candidate offsets $\boldsymbol{\xi}_{ij}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{3})$ in the local frame and map them to world coordinates. To focus sampling on reliable (high-opacity) regions, we apply Bernoulli thinning controlled by $\alpha_{i}$:
@@
We keep $\mathbf{p}_{ij}$ if $b_{ij}=1$, in expectation each Gaussian contributes $\alpha_{i}M$ points when drawing $M$ candidates. Here, $M$ is the maximum number of candidate samples drawn per Gaussian (before opacity-guided thinning) and directly controls the point density and sampling cost. In all experiments we use $M=100$ (i.e., at most 100 candidate draws per splat). Since candidates are retained with probability $\alpha_{i}$, the expected number of retained samples per Gaussian is $\alpha_{i}M$, which keeps the computational budget bounded while still allocating higher sample density to high-opacity (more reliable) regions.
Front-to-back compositing and accumulated alpha. In 3DGS, solidity is not defined by an explicit surface. It is defined implicitly by how Gaussians accumulate opacity under the renderer’s front-to-back alpha compositing. Therefore, any surface sampling rule that aims to match what the model optimized should be expressed using the same compositing mechanism. For a view $v$ and pixel $\mathbf{u}$, the renderer composites Gaussians front-to-back using transmittance tracking. If $\alpha_{k,v}(\mathbf{u})$ is the per-Gaussian alpha contribution at $\mathbf{u}$, we define $T_{0}(\mathbf{u})=1$ and update
@@
The per-pixel accumulated alpha mask $A_{v}(\mathbf{u})$ tells us whether a ray intersects the reconstructed mass somewhere along the ray, but it does not tell us whether a particular 3D sample $\mathbf{p}$ lies on the visible surface, behind it, or in front of it (all can project to the same pixel). For surface extraction we therefore need a depth-aware test that measures how much opacity accumulates in front of $\mathbf{p}$ along the viewing ray, using the same compositing model that governs training. RaDe-GS provides exactly this query by integrating point-wise alphas with a depth-plane formulation.
Depth-aware point-wise opacity integration. Beyond per-pixel alpha, RaDe-GS provides an integration query at an arbitrary 3D point $\mathbf{p}$: it returns the projected pixel coordinate $\mathbf{u}_{v}(\mathbf{p})$ and a cumulative opacity value $\tilde{\alpha}_{v}(\mathbf{p})$ that measures how much opacity is accumulated in front of $\mathbf{p}$ along the viewing ray. Let $r_{v}(\mathbf{p})=\|\mathbf{R}_{v}\mathbf{p}+\mathbf{t}_{v}\|_{2}$ denote the point ray-distance in view $v$. For a Gaussian $k$, RaDe-GS evaluates a ray-space quadratic form using a precomputed inverse covariance over the 3D offset
@@
The $\min(\cdot)$ depth clamping in Equation (6) is critical for surface consistency: once $\mathbf{p}$ lies behind the locally rasterized surface depth $d_{k,v}(\cdot)$, the depth residual term saturates rather than increases with $r_{v}(\mathbf{p})$, preventing spurious “volumetric” accumulation behind the surface. Then, RaDe-GS composites these point-wise alphas using the same front-to-back transmittance update as in Equation (5), yielding the cumulative opacity
@@
Multi-view surface consistency and reliability scores. Since DBH fitting is highly sensitive to even small amounts of geometric contamination, we aggregate evidence across the calibrated views and attach a per-point reliability that downstream steps can threshold or use as a weight. Concretely, we (i) only allow a view to contribute if the point projects to a foreground pixel in that view, and we (ii) summarize the point-wise opacity integrals across contributing views conservatively. Let $\mathcal{P}=\{\mathbf{p}_{ij}\mid b_{ij}=1\}$ be the candidate samples retained after the opacity-guided thinning in Equation (4). For each $\mathbf{p}\in\mathcal{P}$, we evaluate its support over the calibrated views; a view $v$ contributes only if $\mathbf{p}$ projects inside the image and lands on a non-background pixel in the rendered accumulated alpha mask $A_{v}$. Specifically, with $\mathbf{u}_{v}(\mathbf{p})$ and the projection of $\mathbf{p}$, we define
$m_{v}(\mathbf{p})=\mathbf{1}\!\left(A_{v}(\mathbf{u}_{v}(\mathbf{p}))>0\right).$} (9)
@@
Here, $\bar{\alpha}(\mathbf{p})$ is a conservative multi-view opacity estimate, and $w(\mathbf{p})$ counts the number of views in which $\mathbf{p}$ projects onto a foreground pixel, i.e., $A_{v}(\mathbf{u}_{v}(\mathbf{p}))>0$. In practice, for surface extraction, we keep points with $\bar{\alpha}(\mathbf{p})>\tau$ and $w(\mathbf{p})>0$. Especially for this work, we set $\tau=0$ to keep all points while storing $(\bar{\alpha},w)$ for downstream weighting.
@@ 3.5 Semantic Trunk Extraction
The sampled point cloud from Section 3.4 contains trunks mixed with foliage, understory vegetation, and occasional floating samples caused by Gaussians with nonzero opacity in free space. To isolate stem geometry for DBH measurement, we used ForestFormer3D [50], a 3D semantic segmentation model, to predict a class label for each sampled point and to assign per-tree instance IDs when available. In our pipeline, the per-point opacity reliability $\bar{\alpha}$ is kept as an auxiliary output feature alongside $(x,y,z)$. We retain only points predicted as trunk to form a trunk-only cloud for each tree instance $t$, denoted by $\mathcal{T}_{t}=\{(\mathbf{x}_{k},\bar{\alpha}_{k})\}$. This separation step reduces contamination from branches and understory, while the retained opacity values enable reliability-weighted fitting downstream. The trunk-only instances are then passed to the geometric measurement stage described next.
@@ 3.6 Opacity-Weighted DBH Measurement
DBH is defined relative to local ground height. For each tree instance, we estimate a ground elevation $z_{g}$ by querying a digital terrain model (DTM) at the tree location (e.g., nearest-neighbor lookup at the mean trunk $(x,y)$), and we convert all points to height above ground:
@@
We perform all subsequent slicing and DBH evaluation in the $(x,y)$ plane as a function of $h$.
Next, we construct a sequence of slice centers $\{h_{s}\}$ with spacing $\Delta z$ and thickness $H$ (a slab of height $H$), and we collect 2D points in each slab:
@@
Slices with fewer than a small minimum number of points $n_{\min}$ are discarded. This step produces a set of candidate diameters along the lower stem.
Unlike classical TLS stem fitting where points often lie near the circumference, our sampling in Section 3.4 draws points from Gaussian volumes; therefore, many slice points legitimately fall inside the trunk cross-section. For this reason we treat inliers as belonging to a filled disk rather than a narrow ring around a circle. This choice makes the estimator consistent with volumetric sampling and substantially more stable under partial visibility.
Opacity-weighted solid-circle RANSAC in each slice. For slice $s$, we have 2D points $\{(\mathbf{q}_{k},w_{k})\}_{k=1}^{n}$ with $\mathbf{q}_{k}=(x_{k},y_{k})^{\top}$ and opacity weights $w_{k}=\bar{\alpha}_{k}$. Each RANSAC hypothesis performs the following steps:
@@
(ii) Circle fit from three points: Fit the circle through the sampled points using the algebraic form $x^{2}+y^{2}+ax+by+d=0$. We solve for $(a,b,d)$ after subtracting the slice mean (for numerical stability), and we then recover $\mathbf{c}=(-\tfrac{a}{2},-\tfrac{b}{2})^{\top}$ and $r^{2}=\lVert\mathbf{c}\rVert_{2}^{2}-d$ (adding the mean back to $\mathbf{c}$).
@@
Height-wise taper regression and DBH prediction. Even after per-slice robust fitting, some slices can be corrupted by branch attachments, residual foliage, or incomplete sampling. Therefore, we fit a simple taper model on the lower stem using RANSAC with a negative-slope prior:
$\hat{d}(h)=\beta_{0}+\beta_{1}h,\qquad{-\kappa\leq\beta_{1}<0},$ (15)
where RANSAC discards outlier slices using an absolute residual threshold $\epsilon$ on diameter ($\left|\hat{d}_{s}-(\beta_{0}+\beta_{1}h_{s})\right|\leq\epsilon$) and the slope constraint enforces physically plausible taper (diameter should not increase with height over short stem segments). The additional bound $\kappa$ prevents unrealistically steep taper caused by noisy slice diameters. In practice, we fit Equation (15) over progressively larger height ranges starting near the ground (to avoid upper-canopy contamination) until a stable inlier set is found. Finally, we report DBH as:
@@
Role of opacity in robustness. The opacity weights $w_{k}=\bar{\alpha}_{k}$ originate from the multi-view rendering consistency test in Section 3.4. Points that are only weakly supported (e.g., floaters, thin vegetation, or ambiguous geometry) tend to have small $\bar{\alpha}$ and therefore (i) are sampled less often during hypothesis generation and (ii) contribute little to the inlier score in Equation (14). This coupling between rendering consistency and geometric fitting is the central mechanism behind our opacity-weighted measurement mode.